# Separating subsets from their images

### v1

@@ 1. $(a)$
    $\mathrm{Alt}(n)\leq H\leq\mathrm{Sym}(n)$ in their natural action on $n$ points;
@@ 2. $(b)$
    $\mathrm{ASL}_{d}(q)\leq H\leq\mathrm{A}\Gamma\mathrm{L}_{d}(q)$ in their natural action on $q^{d}$ points, where
    $(d,q)\in\{(1,4),(1,5),(1,8),(2,2),(3,2)\};$
  3. $(c)$
    $\mathrm{PSL}_{d}(q)\leq H\leq\mathrm{P}\Gamma\mathrm{L}_{2}(q)$ in their natural action on $\mathrm{PG}_{1}(q)$, where $q\in\{4,5,7,8\}$;
  4. $(d)$
    $H$ is isomorphic to either $M_{11}$ or $M_{12}$ in their natural actions on $11$ or $12$ points, respectively.
Hence, to prove our lemma we need to consider these cases in turn. We will handle cases  $(b),$  $(c)$ and  $(d)$ with $H\cong M_{11}$ by computation at the end of the proof. Now we will provide some useful structural descriptions for cases  $(a)$ and  $(d)$ with $M_{12}$. In case $(a)$, we will use these explicit description to determine which groups meet the upper bound in Lemma 27. For  $(d)$ with $M_{12},$ these will allow us to restrict the search space of the algorithm to a manageable size, so we can handle this but computation at the end of the proof as well.
Let $H$ be a permutation group acting on a finite set $\Omega$, and let $\mathbb{F}_{2}\Omega$ be the permutation module of $H$ over the field $\mathbb{F}_{2}$ with two elements. Note that the wreath product $C_{2}\mathop{\mathrm{wr}}H$ acts imprimitively on $\mathbb{F}_{2}\Omega$. Every imprimitive permutation group $G$ such that a block system $\Sigma$ for $G$ consists of blocks of size $2$, and the permutation group that $G$ induces on $\Sigma$ is isomorphic to $H$, embeds into $C_{2}\mathop{\mathrm{wr}}H$. Moreover, the kernel $K$ of the action of $G$ on $\Sigma$ is an $\mathbb{F}_{2}H$-submodule of $\mathbb{F}_{2}\Omega$, otherwise $K$ would not be normal in $G$. The possibilities for $K$ are, as a consequence, quite limited. Indeed, in both cases, [35] states that $K$ is either trivial, the one-dimensional vector space $D$ spanned by the all-$1$ vector, the (n1)(n-1)-dimensional vector space $A$ containing all the vector whose entries sum equals to zero, and the whole module $\mathbb{F}_{2}\Omega$. We now divide our discussion according to the group $H$.
Suppose that $H=M_{12}$.
  • If $K$ is trivial, $G$ is isomorphic to $H$ as abstract groups. On the other hand, as the blocks of $\Sigma$ have cardinality $2$ and as the stabiliser in $H$ is the stabilizer of a block in $G$, the stabiliser in $G$ has index $2$ in the stabiliser in $H$. As the stabilizer in $H$ is isomorphic to $M_{11}$, which is simple, it contains no subgroups of index $2$. Hence no such permutation group $G$ exists.
  • If $K$ is isomorphic to $D$, $G$ is isomorphic, as abstract groups, to either $C_{2}\times H$ or to the central perfect extension of $H$ by $C_{2}$. In the former case, arguing as before, we found a single possible stabilizer in $G$, which isomorphic to $M_{11}$. In the latter, we observe that, since the Schur multiplier of $M_{11}$ is trivial, the stabilizer of a block of $\Sigma$ is isomorphic to $C_{2}\times M_{11}$. As a consequence, $M_{11}$ is the unique (up to conjugation) subgroup of the central perfect extension, which enjoys the property of being a stabiliser in $G$.
  • If $K$ is isomorphic to $A$, we claim that $G$ is isomorphic, as abstract groups, to the split extension $A\rtimes H$. In fact, if this were not the case, $A\rtimes H$ would have an index $2$ subgroup, which goes against the simplicity of $H$ and the possibilities for $K$. The stabilizer in $G$ is unique (up to conjugation): this follows from the uniqueness of the subgroup $\mathbb{F}_{2}\Omega\rtimes M_{11}$ of index $12$ in the wreath product $\mathbb{F}_{2}\Omega\rtimes M_{12}$, and that, as the action on the blocks is transitive, $H$ acts transitively on the subgroups of index $2$ of $\mathbb{F}_{2}\Omega$. In particular, the stabiliser can be characterized as the intersection of $G$ with the stabilizer of the whole wreath product endowed with the imprimitive action.
  • If $K$ is isomorphic to $\mathbb{F}_{2}\Omega$, the same reasoning as the previous case shows that $G$ coincides with the whole wreath product endowed with the imprimitive action.
Suppose that $H=\mathrm{Alt}(n)$, with $n\geq 8$. (The cases with $n\leq 7$ will be dealt with computationally.) The same reasoning as before can be applied if $K\in\{1,A,\mathbb{F}_{2}\Omega\}$.
  • If $K$ is trivial, we find no possibilities for $G$.
  • If $K$ is isomorphic to $A$, $G$ is the split extension $A\rtimes H$, and the stabilizer is obtained by intersecting $G$ with the stabilizer of the whole wreath product endowed with the imprimitive action.
  • If $K$ is isomorphic to $\mathbb{F}_{2}\Omega$, $G$ is with the whole wreath product endowed with the imprimitive action.
The case $K=D$ is slightly different, as our hypothesis on the degree prescribes that the Schur multipliers of both $\mathrm{Alt}(n)$ and its stabiliser $\mathrm{Alt}(n-1)$ are cyclic of order $2$.
  • If $K$ is isomorphic to $D$, $G$ is isomorphic, as abstract groups, to either $C_{2}\times H$ or to the central perfect extension of $H$ by $C_{2}$. In the former case, as before, $\mathrm{Alt}(n-1)$ is the single possibility for the stabiliser in $G$ (up to conjugation). In the latter, the stabiliser of a block of $\Sigma$ is isomorphic to the central perfect extension of $\mathrm{Alt}(n-1)$ by $C_{2}$. As $\mathrm{Alt}(n-1)$ is simple, this extension does not have any index two subgroup. Hence, if $G$ is the central perfect extension of $H$ by $C_{2}$, $G$ does not admit an imprimitive action of degree $2n$.
Finally, we suppose that $H=\mathrm{Sym}(n)$, with $n\geq 8$. The fact that $H$ contains $\mathrm{Alt}(n)$ as a subgroup of index $2$ makes the treatment of this case more convoluted.
  • If $K$ is trivial, $G$ is isomorphic to $H$ as abstract groups, and the stabilizer in $G$ is the unique (up to conjugation) subgroup of index $2n$ in $\mathrm{Sym}(n)$, that is, $\mathrm{Alt}(n-1)$.
  • If $K$ is isomorphic to $D$, $G$ is isomorphic, as abstract groups, to either $C_{2}\times H$ or to the central perfect extension of $H$ by $C_{2}$. In the former case, as the stabiliser of a block of $\Sigma$ is $C_{2}\times\mathrm{Sym}(n-1)$, there are two subgroups of index $2$ of $C_{2}\times\mathrm{Sym}(n-1)$: it is isomorphic to either $\mathrm{Sym}(n-1)$ or to $\mathrm{Alt}(n-1)\times C_{2}$. Only the first one can be the stabiliser in $G$, because $\mathrm{Alt}(n-1)\times C_{2}$ is not core-free. In the latter, we observe that the stabilizer of a block of $\Sigma$ is isomorphic to the central perfect extension of $\mathrm{Sym}(n-1)$ by $C_{2}$. The unique subgroup of index $2$ of the stabiliser of a block is the central perfect extension of $\mathrm{Alt}(n-1)$ by $C_{2}$, which is not core-free. Hence, $G$ cannot be endowed with an imprimitive action on $2n$ points.
  • If $K$ is isomorphic to $A$, we observe that $A\rtimes\mathrm{Alt}(n)$ has index $4$ as a subgroup of $C_{2}\mathop{\mathrm{wr}}\mathrm{Sym}(n)$, and their quotient is isomorphic to $C_{2}\times C_{2}$. Hence, there are three possibilities for $G$ as an abstract group. If $G$ is isomorphic to the split extension $A\rtimes H$, as we did for $M_{12}$ and $\mathrm{Alt}(n)$, we conclude that the stabilizer in $G$ is isomorphic to the intersection of $G$ with the stabiliser in $C_{2}\mathop{\mathrm{wr}}\mathrm{Sym}(n)$. If $G$ is isomorphic to $C_{2}\mathop{\mathrm{wr}}\mathrm{Alt}(n)$, we note that $G$ induces $\mathrm{Alt}(n)$ as permutation group on $\Sigma$, against our hypothesis. Last, if $s$ denotes an element of the wreath product outside of $A$, the remaining possibility for $G$ is generated as
    $G=\left\langle A,\mathrm{Alt}(n),sg\mid g\in\mathrm{Sym}(n)\setminus\mathrm{Alt}(n)\right\rangle,$
    and its stabilizer is obtained by intersecting $G$ with the whole wreath product.
  • If $K$ is isomorphic to $\mathbb{F}_{2}\Omega$, $G$ coincides with the whole wreath product endowed with the imprimitive action.
We are ready to determine for which groups arising in $(a)$ the upper bound from D is attained.
@@ 4. $(d)$
Let $n\geq 8$ be a positive integer, and let $G$ be an imprimitive group with block system $\Sigma$ that consists of $n$ blocks of size $2$. Suppose that the action of $G$ on $\Sigma$ is permutationally isomorphic to $\mathrm{Alt}(n)$ or $\mathrm{Sym}(n)$. Then $\mathbf{m}(G)=n+1$ if, and only if, $G$ is one of the following:
@@ 1. $(i)$
    $\mathrm{Sym}(n)$ with stabiliser $\mathrm{Alt}(n-1)$ acting on $2n$ points;
@@ 2. $(ii)$
    $C_{2}\times\mathrm{Alt}(n)$;
@@ 4. $(iv)$
    $A\rtimes\mathrm{Alt}(n)$ for even $n$;
@@ 5. $(v)$
    $A\rtimes\mathrm{Sym}(n)$ for even $n$;
@@ 6. $(vi)$
    $\langle A,\mathrm{Alt}(n),sg\mid g\in\mathrm{Sym}(n)\setminus\mathrm{Alt}(n)\rangle$, where $A$, $\mathrm{Alt}(n)$ and $s$ are as before;

### v2

@@ 1. $(a)$
    $\mathrm{Alt}(m)\leq H\leq\mathrm{Sym}(m)$, or
@@ 2. $(b)$
    $m=12$ and $H\cong M_{12}$.
As was proved in [32], the only submodules of the permutational module $\mathbb{F}_{2}^{\Sigma}$ of $H$ are then either the trivial module $\mathbf{1}$, the trace module $\mathsf{T}$, the augmentation module $\mathsf{A}$, and of course $\mathbb{F}_{2}^{\Sigma}$ itself.
Suppose first that the kernel $K$ is trivial. By Lemma 25, the stabiliser of a point in $H$ contains a subgroup $S$ of index $2$, excluding the cases $H\cong\mathrm{Alt}(m)$ and $H\cong M_{12}$. Hence $H=\mathrm{Sym}(m)$ and thus $S=\mathrm{Alt}(m-1)$. Lemma 25 then implies that $G$ is permutation isomorphic to the action of $\mathrm{Sym}(m)$ on the cosets of $\mathrm{Alt}(m-1)$, and thus permutation isomorphic to $\mathop{\mathrm{mix}}(\mathrm{Sym}(m),\mathrm{Alt}(m),\mathbf{1})$, as explained in Remark 26.
Suppose now that $K=\mathsf{T}$. Then $G$ is a central $2$-extension of $H$. Note also that $K=\langle\tau\rangle$ where $\tau$ swaps the two points in every block in $\Sigma$.
If the extension is split with the complement $\overline{H}\cong H$, then $G=C_{2}\times\overline{H}$. Further, if $\overline{H}$ is intransitive, then by Lemma 28, $G=\mathsf{T}\rtimes H$, as in Construction 27. A direct computation in the case of $H=M_{12}$ exhibits a non-self-separable of size $12$ (see [5] for such a set), showing that $\mathbf{m}(\mathsf{T}\rtimes M_{12})\leq 12$, contradicting our assumptions.
On the other hand, if $\overline{H}$ is transitive, then we may apply the previous paragraph with $\overline{H}$ in place of $G$ to conclude that $\overline{H}\cong\mathop{\mathrm{mix}}(\mathrm{Sym}(m),\mathrm{Alt}(m),\mathbf{1})$ in its action on $\Omega^{\prime}:=\{1,\ldots,m\}\times\{0,1\}$. In this setting, the block system $\Sigma$ corresponds to the partition of $\Omega^{\prime}$ into pairs $\{\sigma\}\times\{0,1\}$, and the generator $\tau$ of the kernel $K$ to the permutation $\tau^{\prime}$ of $\Omega^{\prime}$ that swaps both elements in each of these pairs. The permutation group $G$ is thus permutation isomorphic to the action of $\langle\overline{H},\tau^{\prime}\rangle$ on $\Omega^{\prime}$, which is clearly isomorphic to the permutation group $\mathsf{T}\rtimes H$, as in Construction 27. This completes the proof in the case the extension of $K$ by $H$ is split.
Now suppose that $G$ is a non-split extension of $K=\langle\tau\rangle$ by $H$. Let $\pi\colon G\to G^{\Sigma}=H$ be the corresponding epimorphism with kernel $K$. Since the stabiliser $G_{\omega}$ of a point $\omega\in\Omega$ intersects $K$ trivially, $\pi$ maps $G_{\omega}$ isomorphically to the stabiliser $H_{B}$ of a point $B\in\Sigma$. Furthermore, $\pi^{-1}(H_{B})=G_{B}=\langle\tau\rangle\times G_{\omega}$.
Suppose first that $H=\mathrm{Sym}(m)$. Then (as is well known) there exist two non-isomorphic non-split central $2$-extensions, usually denoted by $2.\mathrm{Sym}(m)^{+}$ and $2.\mathrm{Sym}(m)^{-}$. In both cases the central element is contained in the derived subgroup $[P,P]$ of the the preimage $P$ of the group $\mathrm{Sym}(m-1)$. In our context that implies that $\tau\in[G_{B},G_{B}]$, which clearly contradicts the fact that $\tau$ is a direct (central) factor in $G_{B}$. Similarly, if $H=\mathrm{Alt}(m)$, then $G$ is isomorphic to the unique non-split central $2$-extension of $\mathrm{Alt}(m)$, usually denoted by $2.\mathrm{Alt}(m)$. Here it can also be seen that $\tau$ resides within the derived subgroup $G_{B},G_{B}]$, again yielding a contradiction.
We are thus left with the case where $H=M_{12}$. Since $M_{12}$ is a perfect simple group and has the Schur multiplier of order $2$, there exists a unique non-split $2$-extension, namely the group usually denoted by $2.M_{12}$. It can be checked computationally that this group has a unique permutation representation of degree $24$ (admitting blocks of size $2$ and acting on them as $M_{12}$), which can be found in the Database of Transitive Permutation Groups in Magma [9] under the name TransitiveGroup(24,18440)(24,18440). One can check computationally that this group admits a non-self-separable set of size $12$, implying that $\mathbf{m}(G)\leq 12$, contradicting our assumption on $\mathbf{m}(G)$ (see [5] for the computational details). This finishes the proof in the case where $K=\mathsf{T}$.
Let us now assume that $K=\mathsf{A}$. If $H\cong\mathrm{Alt}(m)$ or $M_{12}$, then Lemma 30 implies that $G=\mathsf{A}\rtimes H$. However, if $m$ is odd, then the subset $A:=\{1,\ldots,m\}\times\{0\}$ of the domain of $\mathsf{A}\rtimes H$ (see Construction 27) cannot be separated from itself since $G$ does not contain the trace submodule $\mathsf{T}$. On the other hand, if $H\cong\mathrm{Sym}(m)$, then $G$ is isomorphic either to $\mathsf{A}\rtimes\mathrm{Sym}(m)$ (and as above, $m$ needs to be even) or to $\mathop{\mathrm{mix}}(\mathrm{Sym}(m),\mathrm{Alt}(m),\mathsf{A})$.
Finally, if $K=\mathbb{F}_{2}^{\Sigma}$, then the single option is $G=\mathbb{F}_{2}^{\Sigma}\rtimes H=C_{2}\mathop{\mathrm{wr}}H$. This concludes the proof of the lemma. ∎
This concludes our analysis of permutation groups admitting blocks of size $2$.
7.2.2. Permutation groups admitting two blocks of size $m$.
We start by describing some quirky permutation subgroups of $H\mathop{\mathrm{wr}}C_{2}$.
Construction 33.
Let $H$ be a transitive permutation group of degree $n$, and let $f\in\mathrm{Sym}(n)$ be a permutation such that $f^{2}\in H$ (and thus $f\in\mathbf{N}_{\mathrm{Sym}(n)}(H)$). We denote the permutation group
$\left\{(h,h)\mid h\in H\right\}\cong H,$
whose action is diagonal on $\Omega^{\{0,1\}}$, by $\mathbf{q}_{0}(H)$. Note that $\mathbf{q}_{0}(H)$ is intransitive with two orbits, and that its transitive constituents are permutation isomorphic to $H$. We define
$\mathbf{q}(H,f)=\langle\mathbf{q}_{0}(H),\bar{f}\rangle,$
where
$(\omega,0)^{\bar{f}}=(\omega^{f},1),\quad\hbox{and}\quad(\omega,1)^{\bar{f}}=(\omega^{f},0).$
Note that $\mathbf{q}(H,f)$ is transitive on $\Omega^{\{0,1\}}$, and that $\mathbf{q}_{0}(H)$ is an index $2$ subgroup of $\mathbf{q}(H,f)$. Further, observe that $\mathbf{q}(H,1)=H\times C_{2}$.
33 captures, up to permutational isomorphism, every permutation group which is an imprimitive extension of $H$ of degree $2$, as stated in the following lemma.
@@
Let $H$ be a permutation group on $\Omega$, let $s\in\mathrm{Sym}(\Omega^{\{0,1\}})$ be a permutation with the properties that it swaps the two sets $\Omega^{\{0\}}$ and $\Omega^{\{1\}}$, and that $s^{2}\in\mathbf{q}_{0}(H)$ endowed with its diagonal action on $\Omega^{\{0,1\}}$. Denote by $G$ the permutation group on $\Omega^{\{0,1\}}$ generated by $\mathbf{q}_{0}(H)$ and $s$. Then $G$ is permutation isomorphic to $\mathbf{q}(H,f)$, for some $f\in\mathrm{Sym}(\Omega)$.
Proof.
Aiming for a contradiction, suppose that, for every $f\in\mathrm{Sym}(\Omega)$,
$s\notin{\bar{f}}^{{\bf N}_{\mathrm{Sym}(\Omega^{\{0,1\}}}\left(\mathbf{q}_{0}(H)\right)}.$
In particular, there exist two distinct permutations $a,b\in\mathrm{Sym}(\Omega)$ such that
$(\omega,0)^{s}=(\omega^{a},1),\quad\hbox{and}\quad(\omega,1)^{s}=(\omega^{b},0).$
From the fact that $s^{2}\in\mathbf{q}_{0}(H)$, it follows that
$s\in{\bf N}_{\mathrm{Sym}(\Omega^{\{0,1\}})}\left(\mathbf{q}_{0}(H)\right)=\langle\mathbf{q}_{0}\left({\bf N}_{\mathrm{Sym}(\Omega)}\left(H\right)\right),\bar{1}\rangle.$
But no element in the normaliser has the desired form, which is a contradiction. Hence, $s$ and $\bar{f}$ are conjugate in ${\bf N}_{\mathrm{Sym}(\Omega^{\{0,1\}})}\left(\mathbf{q}_{0}(H)\right)$, which proves the permutational isomorphism (see [17, Exercise 1.6.1]). ∎
Example 35.
We specialize 33 for the alternating and symmetric groups. Let $m\geq 5$ be an integer. Observe that, if $f\in\mathrm{Alt}(m)$ is an even permutation, then $(f,f)\bar{f}=\bar{1}$, and hence
$\mathbf{q}(\mathrm{Alt}(m),f)=\mathbf{q}(\mathrm{Alt}(m),1)=\mathrm{Alt}(m)\times C_{2}.$
Analogously, if $f\in\mathrm{Sym}(m)$ is odd, then, by choosing $g\in\mathrm{Alt}(m)$ such that $gf=(1,2)$, we have $(g,g)\bar{f}={\overline{(1,2)}}$. Therefore,
$\mathbf{q}(\mathrm{Alt}(m),f)=\mathbf{q}(\mathrm{Alt}(m),(1,2)).$
On the other hand, the same reasoning shows that, for every permutation $f\in\mathrm{Sym}(m)$,
$\mathbf{q}(\mathrm{Sym}(m),f)=\mathbf{q}(\mathrm{Sym}(m),1).$
The appearance of such subgroups in our setting is justified by the following observation.
Lemma 36.
Let $H$ be an almost simple primitive permutation group with socle $T$, and let $G$ be an imprimitive permutation group $G$ with two blocks of imprimitivity, $B$ and $C$, such that $G_{B}^{B}=H$. Then, either $T\mathop{\mathrm{wr}}C_{2}\leq G\leq H\mathop{\mathrm{wr}}C_{2}$, or $G$ is permutation isomorphic to $\mathbf{q}(H,f)$, for some $f\in\mathrm{Sym}(\Omega)$ as defined in 33.
Proof.
We start by computing the socle of $G$. We note that $G_{(\Sigma)}=G_{B}=G_{C}$ and that $G$ is an extension of $C_{2}$ by $G_{(\Sigma)}$. Let $N$ be a minimal normal subgroup of $G_{(\Sigma)}$. Consider the projections $\pi_{B}:G_{(\Sigma)}\to G_{B}^{B}$ and $\pi_{C}:G_{(\Sigma)}\to G_{C}^{C}$. Since $N$ is minimal normal, so are $N^{\pi_{B}}$ and $N^{\pi_{C}}$. Observe that, as $G$ is faithful, the two images cannot be concurrently trivial. In particular, $T$ is a subgroup of $N$, and hence the socle of $G$ is of the form $T^{\ell}$, for some positive integer $\ell$. As $G\leq\mathrm{Sym}(|B|)\mathop{\mathrm{wr}}C_{2}$, we see that $\ell\leq 2$, and hence either $\ell=1$ or $\ell=2$. On one hand, if $\ell=2$, it follows at once that
$T\mathop{\mathrm{wr}}C_{2}\leq G\leq H\mathop{\mathrm{wr}}C_{2}.$
On the other hand, suppose that $\ell=1$. Since $G$ acts transitively on $\Sigma$, there exists a permutation $s\in G$ such that $B^{s}=C$ and $s^{2}\in G_{(\Sigma)}$. Observe that $G_{B}^{B}$ and $G_{C}^{C}$ are permutation isomorphic: the requested bijections are $s$ for the domains and the conjugation by $s$ for the groups. By Lemma 34, $G$ is permutation isomorphic to $\mathbf{q}(H,f)$, where $f$ is a permutation with the property that $\bar{f}=s$. ∎
We can now state the main result of the section.
Lemma 37.
Let $m\geq 2$ be a positive integer, and let $G$ be an imprimitive group with block system $\Sigma$ that consists of two blocks of size $m$. Then $\mathbf{m}(G)=m+1$ if and only if $G$ is permutation isomorphic to one of the following:
@@ 1. $(i)$
    $C_{2}\times C_{2}$, $C_{4}$ or $\rm Dih(4)$ acting transitively on $4$ points;
@@ 2. $(ii)$
    $\mathbf{q}(H,1)\cong H\times C_{2}$ (see Construction 33), where $m$ and $H$ satisfy one of:
    • $m\geq 3$ and $H=\mathrm{Sym}(m)$;
    • $m\geq 4$ and $H=\mathrm{Alt}(m)$;
    • $m=4$ and $H=C_{2}\times C_{2}$ or $\rm Dih(4)$;
    • $m=6$ and $H=\rm PGL_{2}(5)$;
@@ 4. $(iv)$
    $H\mathop{\mathrm{wr}}C_{2}\cong H^{2}\rtimes C_{2}$ where $m$ and $H$ satisfy one of:
    • $m\geq 3$ and $H=\mathrm{Sym}(m)$;
    • $m\geq 4$ and $H=\mathrm{Alt}(m)$;
    • $m=4$ and $H=C_{4},\;C_{2}\times C_{2}$ or $\rm Dih(4)$;
    • $m=5$ and $H=\rm AGL_{1}(5)$;
    • $m=6$ and $H=\rm PGL_{2}(5)$;
    • $m=8$ and $H=\rm PSL_{2}(7)$, $\rm PGL_{2}(7)$, $\rm AGL_{1}(8)$, $\rm A\Gamma L(1,8)$ or $\rm ASL_{3}(2)$;
    • $m=9$ and $H=\rm PSL_{2}(8)$ or $\rm P\Gamma L_{2}(8)$;
    • $m=12$ and $H=M_{12}$;
@@ 5. $(v)$
    one of transitive groups of degree $6$ that can be found in the Library of Transitive Groups in [9] under the names TransitiveGroup(6,i)(6,i) for $i\in\{9,10\}$.
@@ 6. $(vi)$
    one of transitive groups of degree $8$ that can be found in the Library of Transitive Groups in [9] under the names TransitiveGroup(8,i)(8,i) for $i\in\{19,21,22,26,28,29,$ 30,31,33,34,41,45,46}30,31,33,34,41,45,46\}.